OpenStudy (anonymous):

complete the statement 15^4a^2+20x^2a^3=5x^2a^2()

OpenStudy (anonymous):

help please!!!!

OpenStudy (anonymous):

@sbuck98

OpenStudy (sbuck98):

I don't know that

OpenStudy (anonymous):

:( math sucks lol

OpenStudy (milkdud0000):

I assumed you missed an x in the first term. 15x^4a^3+20x^2a^3=5x^2a^2 5x²a²(3x²a + 4a) = 5x²a² 3x²a + 4a = 1 a(3x² + 4) = 1 ....1 ----------- = a <==solved for a 3x² + 4 3x²a + 4a = 1 3x²a = 1 - 4a 1 - 4a --------- = x² ...3a ....._____ ..../1 - 4a ../--------- = x <== solved for x \/...3a 15x^4a^3+20x^2a^3=5x^2a^2 15x^4a^3 + 20x^2a^3 - 5x^2a^2 = 0 5x²a²(3x²a + 4a - 1) = 0 <== factored form I didn't know what you wanted, so I tried to cover all the bases. I hope that helps!! :-

OpenStudy (milkdud0000):

the answer is 1/3(1/a-4)